OpenStudy (anonymous):

What are the roots of x in -10x^2 + 12x − 9 = 0?

OpenStudy (anonymous):

@Directrix @MrNebula @ParthKohli

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

so you tried to use quadratic formula?

OpenStudy (freckles):

\[ax^2+bx+c=0 \\ \text{ has solutions } x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

No but I will now.

OpenStudy (mrnebula):

http://xrjunque.nom.es/rootfinder.aspx

OpenStudy (anonymous):